OpenStudy (anonymous):

Please Help Me!!!! What is the solution of the following system 3x+3y=10 -9x-9y=-30

OpenStudy (ranga):

If you multiply the first equation throughout by -3 you get the second equation. So the two equations are basically one and the same equation. There are 2 unknowns and just one equation. That means there are infinite solutions to this problem.

OpenStudy (anonymous):

how u do that

OpenStudy (shamil98):

You could also divide all valuate in the second equation by 3, same result.

OpenStudy (shamil98):

values*

OpenStudy (ranga):

You don't have to solve for x and y. You can just state what I have stated above and say there are infinite number of solutions to this system and leave it at that.

OpenStudy (anonymous):



OpenStudy (anonymous):

need toshow work

OpenStudy (kainui):

@superraymond you're dividing, not multiplying.

OpenStudy (shamil98):

I said dividing the second equation not the first, if you are using my method.

OpenStudy (anonymous):

im confuse

OpenStudy (raffle_snaffle):

system of equations!

OpenStudy (anonymous):

yea im so lost

OpenStudy (kainui):

So some tips about multiplying equations. First thing, the equals sign literally means both sides have the same thing on both sides, even if they look different, they are exactly the same. So if I say I have an equal number of pebbles in both of my hands, that means if I add a pebble to one hand, I need to add a pebble to the other hand for them to remain equal. I can also double (multiply by 2) the amount of rocks in one hand and since they're equal if I double the amount in the other hand, then they must still be equal! Now suppose I tell you I have 3 rocks plus a secret number more of rocks all in my left hand then I say I have 5 rocks in my right hand and say that I have an equal number of rocks in both hands, that means that secret number must be 2 rocks. So an equation would look like: 3+x=5 Hopefully this helps. Ask more questions if you need more help! =)

OpenStudy (anonymous):

can u show me how to slove it

OpenStudy (kainui):

I can't/won't. Someone else might.

OpenStudy (anonymous):

why not

OpenStudy (mathmale):

SuperRaymond: In discussing solutions to systems of two linear equations (such as you have to deal with here), there are three cases to be learned: Case I: When drawn on the same set of axes, the graphs of the two lines cross in just one place: Here you have ONE unique solution to the system of linear equations. Case 2: When drawn on the same set of axes, the graphs of the two lines never cross; in other words, they're parallel to each other: Here you have NO solution. Case 3: When drawn on the same set of axes, the graphs of the two lines are one and the same; they touch at infinitely many places along their lengths: Here you have infinitely many solutions. In the problem you've presented, Case 3 applies.

OpenStudy (anonymous):

but I must show my work how to do that

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

From 2nd equation get x or y. then put in 1st equation

OpenStudy (mathmale):

Let's go through that work. You have the system of linear equations 3x+3y=10 -9x-9y=-30 If you multiply every term of the first equation by -3, the result will be exactly the same as the second equation (right?) Thus, the two lines are one and the same; the two equations are actually one and the same. This is Case 3, as explained earlier. This is all the work that you need to show. Again, Super Raymond, you must learn those three cases and be able to identify which you have in a given problem involving systems of linear equations.

OpenStudy (anonymous):

how u get -3

OpenStudy (mathmale):

Super R: Stop and think for a moment. Where do you think I got that -3, and why do you think I wanted to multiply the first equation by -3?

OpenStudy (anonymous):